# IMPROVE: Iterative Model Pipeline Refinement and Optimization Leveraging LLM Agents

### v1

@@ Abstract
Computer vision is a critical component in a wide range of real-world applications, including plant monitoring in agriculture and handwriting classification in digital systems. However, developing high-performance computer vision models traditionally demands both machine learning (ML) expertise and domain-specific knowledge, making the process costly, labor-intensive, and inaccessible to many. Large language model (LLM) agents have emerged as a promising solution to automate this workflow, but most existing methods share a common limitation: they attempt to optimize entire pipelines in a single step before evaluation, making it difficult to attribute improvements to specific changes. This lack of granularity leads to unstable optimization and slower convergence, limiting their effectiveness. To address this, we introduce Iterative Refinement, a novel strategy for LLM-driven ML pipeline design inspired by how human ML experts iteratively refine models, focusing on one component at a time rather than making sweeping changes all at once. By systematically updating individual components based on real training feedback, Iterative Refinement improves stability, interpretability, and overall model performance. We implement this strategy in IMPROVE, an end-to-end LLM agent framework for automating and optimizing object classification pipelines. Through extensive evaluations across datasets of varying sizes and domains, including standard benchmarks and Kaggle competition datasets, we demonstrate that Iterative Refinement enables IMPROVE to consistently achieve better performance over existing zero-shot LLM-based approaches. These findings establish Iterative Refinement as an effective new strategy for LLM-driven ML automation and position IMPROVE as an accessible solution for building high-quality computer vision models without requiring ML expertise.
@@ 1 Introduction
Computer vision has emerged as a powerful tool for addressing many complex real-world problems, from plant monitoring in agriculture to handwriting classification in digital systems. However, the process of training computer vision models has grown increasingly complex, involving many different steps such as data augmentation, architecture selection, and hyperparameter tuning. As a result, developing high-performing models is labor-intensive and often demands machine learning (ML) expertise, as well as domain-specific knowledge. Each component must be manually calibrated based on prior experience and detailed analysis of training statistics.
To simplify this process, tools such as Weights & Biases have been developed, offering ML practitioners the ability to track, organize, and optimize model training workflows. Similarly, researchers have been investigating approaches, such as hyperparameter optimization (Bergstra and Bengio, 2012; Bergstra et al., 2011; Snoek et al., 2012; Springenberg et al., 2016; Falkner et al., 2018) and neural architecture search (Elsken et al., 2017; Kandasamy et al., 2018), to automate parts of the ML pipeline. While these platforms and methods provide valuable support, they tend to address isolated aspects of the model development process and still require substantial involvement from human experts.
Recent advances in large language models (LLMs) have opened up new possibilities for automation by enabling AI agents to generate code and reason about design choices. Inspired by this capability, researchers have begun leveraging LLM agents to automate the construction and optimization of ML models (Grosnit et al., 2024; Trirat et al., 2024; Hong et al., 2024a; Li et al., 2024; Guo et al., 2024). Approaches in this emerging field vary widely in terms of planning, reasoning, and evaluation strategies. However, one common feature of most methods is to optimize the entire pipeline simultaneously before each evaluation attempt, rather than refining each component in isolation. While this "all-at-once" approach can help with rapid early-stage improvements, our experiments reveal a key limitation: as model performance approaches its maximum potential, this strategy struggles to make further progress. The inability to attribute performance changes to specific components makes it difficult to refine the pipeline and push improvements beyond a certain threshold.
To address these limitations, we introduce Iterative Refinement, a new strategy for LLM-driven ML system design inspired by how human experts approach model development. Rather than attempting to optimize an entire pipeline at once, experienced ML practitioners typically analyze performance, adjust specific components, and iteratively refine the design based on training feedback. For example, if they determine that data augmentation needs improvement, they will experiment with different augmentation techniques while keeping the training procedure and model architecture unchanged. Only after evaluating its impact will they refine other aspects of the pipeline. Following this intuition, Iterative Refinement systematically updates one component at a time, evaluates its impact, and refines further using real training feedback. This structured approach enables more stable, interpretable, and controlled improvements by isolating the effects of each change and precisely identifying what drives performance gains.
We implement this strategy in IMPROVE (Iterative Model Pipeline Refinement and Optimization leveraging LLM agents), a fully autonomous framework for designing and optimizing image classification pipelines. IMPROVE emulates the workflow of a team of human ML engineers, taking a dataset as input and autonomously constructing, training, and iteratively refining a model through its multi-agent system. It efficiently manages key tasks such as data preprocessing, architecture selection, and hyperparameter tuning.
Our experiments demonstrate that Iterative Refinement enables IMPROVE to generate higher-performing models consistently compared to modifying entire pipelines at once. IMPROVE also greatly outperforms zero-shot LLM-generated training pipelines and achieves near-human-level performance on standard datasets, including CIFAR-10, TinyImageNet, and various Kaggle competition datasets. Notably, it maintains computational efficiency comparable to that of human practitioners, making it both effective and scalable. Our findings validate Iterative Refinement as a practical strategy for LLM-driven ML automation but also establish IMPROVE as an accessible tool for developing state-of-the-art image classification models without requiring deep ML expertise.
Overall, our work introduces a novel strategy for LLM-based ML system design along with a concrete framework that automates the generation and optimization of object classification pipelines. Specifically, our contributions are:
  • We introduce Iterative Refinement, a new design approach in LLM agent frameworks inspired by how human experts iteratively refine models, optimizing individual components incrementally for more stable and controlled improvements.
  • We design IMPROVE, a novel framework that applies Iterative Refinement to create and optimize object classification pipelines using specialized LLM agents.
  • Through extensive experiments on standard and real-world datasets, we demonstrate that Iterative Refinement ensures more consistent performance gains in pipeline optimization, helping IMPROVE to outperform zero-shot LLM-generated pipelines and achieves human expert-level results.
@@ 2 Related Works
2.1 Large Language Models and LLM Agents
Large language models (LLMs) are pre-trained on massive text corpora, often with millions to trillions of parameters. Some of the most well-known LLMs today include GPT-3.5 and GPT-4 (OpenAI et al., 2024) from OpenAI, Claude 3.5 from Anthropic, Mixtral from Mistral, and Llama 3 from Meta. While LLMs can be used directly after pre-training, they often undergo additional training stages, such as supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022). These stages aim to align LLMs with specific behaviors and objectives, resulting in models capable of performing tasks such as general-purpose chatbots, code generation, and information retrieval.
A subset of LLMs specializes in code generation. Codex (Chen et al., 2021), for example, is a GPT model fine-tuned on publicly available GitHub code with a focus on Python programming. Following Codex, Code Llama was introduced, extending coding capabilities to various programming languages such as Python, C++, Java, PHP, and TypeScript. More recently, larger natural language focused LLMs, including GPT-3.5 and Llama 3-8B, have also demonstrated significant coding proficiency across multiple languages, without needing to be explicitly fine-tuned on code.
LLMs have also been used to create human-like agents capable of executing complex instructions autonomously. Examples of such LLM-based agents include ReAct (Yao et al., 2023), Reflexion (Shinn et al., 2023), and SwiftSage Lin et al. (2023), which have demonstrated effectiveness in complex reasoning and decision-making tasks. Expanding on this, works like AutoGPT (Significant Gravitas, 2024) enabled LLM agents to move beyond reasoning and autonomously perform actions, such as executing code and receiving feedback from outputs. Another line of research explores the collaboration of multiple LLM agents, each with specialized roles (Hong et al., 2024b; Du et al., 2024; Park et al., 2023). These multi-agent frameworks have shown that LLMs can effectively assume distinct roles (Tseng et al., 2024), and role specialization enhances their ability to retrieve relevant knowledge, outperforming single-agent systems on tasks requiring reasoning and strategic planning (Sreedhar and Chilton, 2024).
2.2 Automated Machine Learning
Automated Machine Learning (AutoML) is a field of research seeking to automate various stages of the ML pipeline, making it more efficient to develop ML models. Most works focus on automating specific parts of the process, such as data preparation, model architecture selection, and hyperparameter optimization (He et al., 2021).
For instance, works like AutoAugment (Cubuk et al., 2019), Faster AutoAugment (Hataya et al., 2019), DADA (Li et al., 2020), and TrivialAugment (Müller and Hutter, 2021) concentrate on automating the data augmentation component of image classification by applying search strategies over predefined search spaces. Another major focus in AutoML is neural architecture search (NAS), which aims to find the most optimal model architecture from a given search space. A prominent example is Elsken et al. (2017), where a hill-climbing algorithm is used to identify the best convolutional neural network (CNN) architecture. More advanced approaches, such as those by Kandasamy et al. (2018), leverage Bayesian optimization and optimal transport to refine this search.
In addition to NAS, hyperparameter optimization (HPO) is a closely related area that seeks to identify the best set of hyperparameters for a fixed architecture. Random search (Bergstra and Bengio, 2012) is a simple yet efficient method that outperforms traditional grid search (Bergstra et al., 2011) by exploring the search space more effectively. To further enhance the search process, many works incorporate Bayesian optimization in HPO (Snoek et al., 2012; Springenberg et al., 2016; Falkner et al., 2018), enabling better use of past information to guide future decisions.
2.3 AutoML with LLMs
With the recent advancement of LLMs, many researchers have started exploring using LLMs to tackle problems in AutoML, as LLMs offer much greater flexibility in the search space over traditional methods. One of the earliest works in this direction is Yu et al. (2023), which utilizes a fine-tuned Generative Pre-Trained Transformer (GPT) model to create new architectures using crossover, mutation, and selection strategies. Later, many works took advantage of the pre-training of LLMs, which grants them an immense amount of implicit knowledge on neural architectures. GENIUS (Zheng et al., 2023) is one such work, where NAS is performed by simply prompting GPT-4 to generate different configurations, which are evaluated iteratively for a pre-determined number of iterations. In a different direction, Hollmann et al. (2024) proposed the CAAFE method that uses LLMs to automate feature engineering by generating semantically meaningful features for tabular datasets. Zhang et al. (2023) worked on a more comprehensive pipeline, attempting to simultaneously tackle many tasks including object detection, object classification and question answering with LLM-based AutoML, by optimizing the pipeline with LLM-predicted training log. Another work, AutoMMLab (Yang et al., 2024), focuses on automating computer vision model training, guiding LLMs to handle data selection, model selection, and hyperparameter tuning sequentially, based on natural language requests.
Recently, researchers have advanced the integration of LLM agents into AutoML, evolving their use from simple code generation tools to comprehensive systems that can automate a wider range of components in the machine learning pipeline. For instance, Agent K (Grosnit et al., 2024) proposes an end-to-end pipeline, where the agent scrapes data from a Kaggle URL, trains a model, and submits results, leveraging LLM reasoning and Bayesian optimization in the process to optimize the pipeline. Similarly, AutoML-Agent (Trirat et al., 2024) employs a multi-agent framework that automates model generation in a multi-stage pipeline, such as data preprocessing, model design, and hyperparameter optimization, and includes multi-stage verification to ensure correctness. Data Interpreter (Hong et al., 2024a) approaches AutoML through dynamic task decomposition, constructing, adjusting, and executing task and action graphs where each node represents a specific sub-task. Other frameworks focus specifically on data science. For example, AutoKaggle (Li et al., 2024) targets tabular data with a similar multi-agent setup, but includes a library of machine learning functions like missing value imputation and one-hot encoding for the agents to use. DS-Agent(Guo et al., 2024) develops models for text, time series, and tabular data by utilizing case-based reasoning with insights from Kaggle technical reports and code. Although many works in this direction aim to maximize automation, some others focus on human-AI collaboration. For instance, LAMBDA (Sun et al., 2024) is a human-in-the-loop data science agent that employs multi-agent collaboration for code generation, verification, and execution, similar to other methods, but is able to assist users interactively through an intuitive interface.
@@ 3 Method
3.1 Overview
Traditional machine learning (ML) pipeline optimization requires deep ML and domain expertise, as practitioners iteratively adjust data processing, model selection, and training configurations based on training performance. While effective, this traditional workflow is labor-intensive, requiring constant supervision from ML experts at every stage of development.
Recent LLM-based automation methods (Grosnit et al., 2024; Trirat et al., 2024; Hong et al., 2024a; Li et al., 2024; Guo et al., 2024) aim to replace this process by autonomously generating ML pipelines. These methods typically incorporate data-driven reasoning and employ optimization techniques to iteratively refine the pipeline for improved performance. However, most existing approaches optimize the entire pipeline as a single entity and treat performance improvements globally all at once, diverging from human workflows, where ML practitioners refine individual components in an iterative manner.
Global optimization strategies present a number of fundamental challenges that hinder its effectiveness. First, modifying multiple pipeline components simultaneously reduces interpretability, making it difficult to attribute performance changes to specific adjustments. Second, due to this lack of interpretability and the vast search space, performance improvements often occur through chance-based trial and error rather than systematic refinement, leading to instability in results. Finally, while these strategies may initially yield rapid improvements in early iterations, they tend to plateau as they struggle to make meaningful refinements at later stages when the overall performance is already high, resulting in an overall slow convergence.
3.2 Iterative Refinement
To address these limitations, we introduce Iterative Refinement, a structured optimization strategy inspired by how human ML practitioners systematically adjust individual pipeline components based on empirical feedback. Applying Iterative Refinement in an LLM agent framework requires dividing the target pipeline into distinct components based on the problem domain. Each component must be generated independently yet easily integrated into a complete pipeline. Following the initial training and evaluation of the pipeline, one single component is selected for modification in an attempt to improve while the remaining components remain unchanged. The updated pipeline undergoes training and evaluation, and modifications are retained only if they result in performance improvements. Otherwise, the system reverts to the previous configuration. The process repeats for a fixed number of times, allowing the framework to refine the pipeline progressively.
Compared to existing methods that apply holistic optimization methods, Iterative Refinement offers three key advantages:
  • Improved Interpretability: By modifying only one component at a time, performance changes can be directly attributed to specific adjustments.
  • Stable Improvement: The framework does not rely on finding great combinations of parameters and methods based on chance. Instead, it achieves gradual and systematic improvement across the given iterations.
  • Accelerated Convergence: Targeted refinements prevent redundant modifications and ensure a structured optimization trajectory, leading to faster and more efficient performance gains.

### v2

# IMPROVE: Iterative Model Pipeline Refinement and Optimization Leveraging LLM Experts

@@ Abstract
Large language model (LLM) agents have emerged as a promising solution to automate the workflow of machine learning, but most existing methods share a common limitation: they attempt to optimize entire pipelines in a single step before evaluation, making it difficult to attribute improvements to specific changes. This lack of granularity leads to unstable optimization and slower convergence, limiting their effectiveness. To address this, we introduce Iterative Refinement, a novel strategy for LLM-driven ML pipeline design inspired by how human ML experts iteratively refine models, focusing on one component at a time rather than making sweeping changes all at once. By systematically updating individual components based on real training feedback, Iterative Refinement improves overall model performance. We also provide some theoretical edvience of the superior properties of this Iterative Refinement. Further, we implement this strategy in IMPROVE, an end-to-end LLM agent framework for automating and optimizing object classification pipelines. Through extensive evaluations across datasets of varying sizes and domains, we demonstrate that Iterative Refinement enables IMPROVE to consistently achieve better performance over existing zero-shot LLM-based approaches.
@@ 1 Introduction
Designing high-performing machine learning (ML) pipelines is a complex, multi-step process involving data preprocessing, model selection, and training configuration. Despite the availability of tools such as Weights & Biases and progress in hyperparameter optimization (Bergstra and Bengio, 2012; Bergstra et al., 2011; Snoek et al., 2012; Springenberg et al., 2016; Falkner et al., 2018) and neural architecture search (Elsken et al., 2017; Kandasamy et al., 2018), most automated solutions optimize individual components in isolation and still require expert intervention.
Recent advances in large language models (LLMs) have made it feasible to automate ML pipeline construction end-to-end (Grosnit et al., 2024; Trirat et al., 2024; Hong et al., 2024a; Li et al., 2024; Guo et al., 2024). These LLM agent frameworks generate training code, reason about design choices, and coordinate pipeline assembly. However, most current systems attempt to optimize all pipeline components simultaneously. This “global update” strategy suffers from three drawbacks: difficulty attributing improvements to specific components, lack of stability due to entangled changes, and slower convergence as performance saturates.
We address these limitations with Iterative Refinement, a structured strategy that updates one pipeline component at a time. Inspired by how human practitioners work, our approach evaluates the impact of each change in isolation and only accepts modifications that improve performance. We implement this strategy in IMPROVE (Iterative Model Pipeline Refinement and Optimization leVeraging LLM Experts), a multi-agent framework that autonomously builds, evaluates, and refines image classification pipelines.
Our contributions are both empirical and theoretical:
  • We propose Iterative Refinement, a novel agent coordination strategy for LLM-driven ML systems that improves stability, interpretability, and performance attribution.
  • We develop IMPROVE, a fully autonomous multi-agent system that applies Iterative Refinement to optimize pipelines across data preprocessing, model architecture, and training configuration.
  • We establish theoretical guarantees showing that Iterative Refinement offers faster convergence and greater cost-efficiency than global updates under realistic assumptions about search space structure and LLM resource constraints.
  • We validate IMPROVE across standard benchmarks and real-world Kaggle datasets, demonstrating that it achieves human-competitive performance and consistently outperforms global and zero-shot baselines.
@@ 2 Related Works
Large Language Models and LLM Agents. Large language models (LLMs), such as GPT-3.5 and GPT-4 (OpenAI et al., 2024), Claude 3.5, Mixtral, and Llama 3, are pre-trained on large-scale corpora and aligned via supervised fine-tuning and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), enabling capabilities in general-purpose reasoning, code generation, and retrieval.
Code-oriented LLMs like Codex (Chen et al., 2021) and Code Llama extend support for multiple programming languages. Recent general-purpose LLMs also exhibit strong coding abilities without code-specific fine-tuning.
LLMs have been used to build autonomous agents for reasoning and acting, such as ReAct (Yao et al., 2023), Reflexion (Shinn et al., 2023), SwiftSage (Lin et al., 2023), and AutoGPT (Significant Gravitas, 2024). Multi-agent frameworks (Hong et al., 2024b; Du et al., 2024; Park et al., 2023) further enhance capabilities by assigning specialized roles (Tseng et al., 2024), which improves knowledge retrieval and planning (Sreedhar and Chilton, 2024).
Automated Machine Learning. AutoML aims to automate components of the ML pipeline such as data preparation, model selection, and hyperparameter tuning (He et al., 2021). Many efforts focus on individual tasks: AutoAugment (Cubuk et al., 2019), Faster AutoAugment (Hataya et al., 2019), DADA (Li et al., 2020), and TrivialAugment (Müller and Hutter, 2021) automate data augmentation via search over predefined policies.
Neural architecture search (NAS) identifies optimal model structures. Early methods used heuristic search (Elsken et al., 2017), while more recent work incorporates Bayesian optimization and optimal transport (Kandasamy et al., 2018). Similarly, hyperparameter optimization (HPO) improves training configurations: random search (Bergstra and Bengio, 2012) surpasses grid search (Bergstra et al., 2011), and Bayesian strategies (Snoek et al., 2012; Springenberg et al., 2016; Falkner et al., 2018) further refine this process by leveraging performance history.
AutoML with LLMs. LLMs offer a flexible search space for AutoML, inspiring a growing body of work. Early methods such as GPT-NAS (Yu et al., 2023) used fine-tuned LLMs for neural architecture search (NAS) via crossover and mutation strategies. Later, GENIUS (Zheng et al., 2023) prompted GPT-4 directly for NAS configurations. Beyond architectures, CAAFE (Hollmann et al., 2024) automated feature engineering for tabular data, while AutoML-GPT (Zhang et al., 2023) tackled end-to-end vision and NLP tasks. AutoMMLab (Yang et al., 2024) enabled natural language–driven training pipeline generation for computer vision.
Recent efforts expand LLMs into full agent systems. Agent K (Grosnit et al., 2024) automates Kaggle workflows with LLM planning and Bayesian optimization. AutoML-Agent (Trirat et al., 2024) and Data Interpreter (Hong et al., 2024a) use multi-agent coordination for pipeline refinement via staged reasoning and graph execution. Other systems target domain-specific automation: AutoKaggle (Li et al., 2024) and DS-Agent (Guo et al., 2024) handle tabular, text, and time series data through libraries and case-based reasoning. LAMBDA (Sun et al., 2024) enables human-in-the-loop collaboration with agent-driven code generation and execution.
@@ 3 Method
Most LLM-based AutoML systems attempt to optimize the entire pipeline in one step, but this global strategy leads to entangled updates, unstable convergence, and poor attribution of improvement. In contrast, we advocate for an iterative, component-wise refinement strategy that mirrors how human experts systematically tune ML pipelines.
We begin by formalizing this approach and proving theoretical guarantees for its convergence and efficiency. We then introduce the IMPROVE framework, a multi-agent system that implements this strategy, and describe key engineering mechanisms that enable robust execution.
3.1 Iterative Refinement and Its Theoretical Guarantee
We formalize the ML pipeline as a composition of $k$ sequential components:
$F^{(t)}=S_{k}^{(t)}\circ S_{k-1}^{(t)}\circ\cdots\circ S_{1}^{(t)},$ (1)
where $S_{i}^{(t)}\in\mathcal{S}_{i}$ denotes the configuration of the $i$-th module (e.g., data augmentation, model architecture, training schedule) at iteration $t$. Given a dataset $D$ and loss function $\mathcal{L}$, the optimization objective is to minimize:
$\mathcal{L}(F^{(t)}(D))=\mathcal{L}(S_{k}^{(t)}\circ\cdots\circ S_{1}^{(t)}(D)).$
While prior LLM-based systems attempt to jointly optimize all modules at once, effectively updating the entire tuple $(S_{1}^{(t)},\ldots,S_{k}^{(t)})$, this global strategy suffers from high-dimensional search complexity and poor attribution of performance changes. In contrast, Iterative Refinement updates only one module $S_{j}$ at a time:
$F^{(t+1)}=S_{k}^{(t)}\circ\cdots\circ S_{j}^{(t+1)}\circ\cdots\circ S_{1}^{(t)},$
where $S_{j}^{(t+1)}$ is chosen to minimize loss while holding all other components fixed.
This strategy ensures that each update is:
  • Localized: The impact of each change is directly measurable under the current pipeline context.
  • Monotonic: By accepting only improvements, we enforce $\mathcal{L}(F^{(t+1)}(D))\leq\mathcal{L}(F^{(t)}(D))$.
  • Efficient: The effective search space per iteration is restricted to a single module, reducing variance and computational cost.
We now analyze the convergence and efficiency properties of Iterative Refinement (IR) under a formal optimization framework. The pipeline is modeled as Eq. 1. The goal is to minimize the loss $\mathcal{L}(F^{(t)}(D))$ on a dataset $D$.
Assumptions. Our analysis relies on the following assumptions, which reflect realistic properties of LLM-driven AutoML systems:
  • A1. Modular Decomposability: The pipeline consists of $k$ sequential, reconfigurable modules. Each module’s change impacts downstream modules but does not retroactively affect upstream ones. This follows from the ordered, feedforward nature of most ML pipelines.
  • A2. Per-Iteration Candidate Set: At each iteration $t$, the system generates candidate configurations $\mathcal{C}_{j}^{(t)}\subset\mathcal{S}_{j}$ for each module $j$. For IR, only one module is updated at a time from its candidate set; for global optimization, the full tuple is sampled from $\mathcal{C}_{\text{global}}^{(t)}\subset\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{k}$.
  • A3. Improvement Decay (LLM Saturation): The expected gain from global (full-pipeline) optimization decays geometrically over time:
    $\Delta_{\text{global}}^{(t)}\leq\alpha\cdot\gamma^{t},\quad\text{for }0<\gamma<1.$
    This reflects the empirical observation that LLM-generated full-pipeline proposals degrade in effectiveness over successive rounds of self-editing. As the pipeline improves, the probability that random full-scale edits yield further gains diminishes. Additionally, LLMs tend to repeat or overwrite past decisions without introducing truly novel configurations (Shinn et al., 2023), especially in long-horizon tasks. This leads to performance saturation and diminishing returns. Similar trends have been reported in recent agent-based systems (Grosnit et al., 2024; Hong et al., 2024a), where later rounds often plateau unless guided by finer-grained updates.
  • A4. Local Update Stability: The improvement from component-wise updates is bounded below:
    $\Delta_{\text{local}}^{(t)}\geq\beta>0.$
    This models the practical reliability of targeted refinements: when a single module is modified and evaluated in context, the feedback is clearer and more attributable, enabling more effective learning. Empirically, iterative single-component changes have been shown to maintain stable improvement over longer optimization trajectories (Lin et al., 2023; Trirat et al., 2024). Furthermore, because local edits preserve the rest of the pipeline, they reduce the risk of regression and allow the system to exploit the current solution state effectively—especially when downstream agents can build on past context without unexpected interference.
Main Result. We now establish the convergence advantage of Iterative Refinement over global optimization as below.
Theorem 3.1 (Crossover Point for Local Dominance)
Let $f(T)$ denote the cumulative performance gap between local and global strategies after $T$ iterations:
$f(T):=\beta\cdot T-\frac{\alpha\cdot\gamma(1-\gamma^{T})}{1-\gamma}.$
Then $f(T)$ is strictly increasing, and there exists a crossover point
$T^{*}:=\min\{T\in\mathbb{N}\mid f(T)\geq 0\}$
after which local optimization outperforms global optimization in cumulative gain:
$\sum_{t=1}^{T}\Delta_{\text{local}}^{(t)}\geq\sum_{t=1}^{T}\Delta_{\text{global}}^{(t)}\quad\text{for all }T\geq T^{*}.$
Resource-Constrained Optimization.
While convergence rate is critical, real-world LLM-based systems must also operate under strict resource budgets, especially when implemented through multi-agent prompting architectures. Each optimization step incurs a nontrivial cost due to prompt construction, inter-agent communication, and accumulated context length.
To compare global and local optimization in this setting, we analyze their token-level cost structure. Specifically, we decompose the total resource consumption into three components:
  • $c_{0}$: the fixed token cost for system initialization, including loading LLM agents, setting global task definitions, and embedding shared context;
  • $c_{1}$: the per-step cost of updating a single module (e.g., composing prompts and interpreting responses);
  • $c_{2}$: the amortized cost of logging update history per module per step (e.g., appending "agent $i$ modified component $j$ with config $x$").
Under this formulation, we quantify the cumulative token cost of performing $k$ optimization iterations under two strategies:
$g_{\text{global}}(k)=c_{0}+pc_{1}+pc_{2}(k-1),\qquad g_{\text{local}}(k)=c_{0}+c_{1}+c_{2}(k-1),$
where $p$ is the number of modules in the pipeline. Intuitively, global updates modify all $p$ components per step and log full-context history, while local updates affect only one module at a time, leading to more efficient amortization over time.
To compare convergence under equal-cost budgets, we express the ratio of effective update counts (i.e., the number of steps that can be executed within the same token budget):
$\frac{N_{\text{global}}}{N_{\text{local}}}=\frac{g_{\text{global}}(k)}{g_{\text{local}}(k)}=\frac{c_{0}+px}{c_{0}+x},\quad\text{where }x:=c_{1}+c_{2}(k-1).$
We now ask: given this cost gap, can local updates still outperform global updates under a fixed resource budget?
To assess convergence under a fixed resource budget, we model the optimization objective as a smooth and strongly convex function $f(x)$, where $x\in\mathbb{R}^{d}$ denotes the full pipeline configuration. We start with the following assumptions.
  • $f$ is $\mu$-strongly convex: for all $x,y$, we have $f(y)\geq f(x)+\nabla f(x)^{\top}(y-x)+\frac{\mu}{2}\|y-x\|^{2};$
  • $f$ is $L$-smooth: its gradient is Lipschitz continuous with constant $L$, i.e., $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|;$
  • For coordinate-wise updates, the gradient is coordinate-wise Lipschitz with constant $L_{\max}$, i.e., each partial derivative $\nabla_{i}f$ satisfies: $|\nabla_{i}f(x+\delta e_{i})-\nabla_{i}f(x)|\leq L_{\max}|\delta|.$
Under these assumptions, global (full-pipeline) updates using gradient descent satisfy:
$f(x^{(T)})-f(x^{*})\leq\left(1-\frac{\mu}{L}\right)^{T}\cdot\left(f(x^{(0)})-f(x^{*})\right),$
while component-wise updates using Gauss-Southwell coordinate descent satisfy:
$f(x^{(T)})-f(x^{*})\leq\left(1-\frac{\mu_{1}}{L_{\max}}\right)^{T}\cdot\left(f(x^{(0)})-f(x^{*})\right),$
where $\mu_{1}\in\left[\frac{\mu}{p},\mu\right]$ is the effective strong convexity along coordinate directions (see Richtárik and Takáč (2014)).
Lemma 3.2 (Cost-Normalized Superiority Condition)
Under fixed budget, iterative refinement outperforms global updates if:
$\frac{c_{0}+x}{c_{0}+px}<\log_{1-\mu/L}\left(1-\frac{\mu_{1}}{L_{\max}}\right).$
Interpretation and Practical Implications.
Theorem 3.1 formalizes a phenomenon often observed in practice: global updates deliver quick early gains but rapidly stagnate due to entangled edits, saturation, and model instability. In contrast, Iterative Refinement maintains steady progress, making it the dominant strategy beyond a modest number of iterations. The threshold $T^{*}$ quantifies this crossover and is empirically small (e.g., $T^{*}\approx 2.5$ with $\alpha=0.3$, $\beta=0.1$, $\gamma=0.5$).
Lemma 3.2 shows that even when both strategies are constrained to the same token budget (a key constraint in LLM deployment), IR achieves better convergence due to lower per-step overhead and better amortization of memory costs. This advantage is magnified in highly modular systems (large $p$) or when execution spans many steps (large $k$), making IR especially well-suited for multi-agent systems like IMPROVE, which will be introduced below.